# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v3

@@ Abstract
The effects of ligand binding on protein structures and their in vivo functions carry numerous implications for modern biomedical research and biotechnology development efforts such as drug discovery. Although several deep learning (DL) methods and benchmarks designed for protein-ligand docking have recently been introduced, to date no prior works have systematically studied the behavior of docking methods within the practical context of (1) using predicted (apo) protein structures for docking (e.g., for broad applicability); (2) docking multiple ligands concurrently to a given target protein (e.g., for enzyme design); and (3) having no prior knowledge of binding pockets (e.g., for pocket generalization). To enable a deeper understanding of docking methods’ real-world utility, we introduce PoseBench, the first comprehensive benchmark for practical protein-ligand docking. PoseBench enables researchers to rigorously and systematically evaluate DL docking methods for apo-to-holo protein-ligand docking and protein-ligand structure generation using both single and multi-ligand benchmark datasets, the latter of which we introduce for the first time to the DL community. Empirically, using PoseBench, we find that all recent DL docking methods but one fail to generalize to multi-ligand protein targets and also that template-based docking algorithms perform equally well or better for multi-ligand docking as recent single-ligand DL docking methods, suggesting areas of improvement for future work. Code, data, tutorials, and benchmark results are available at https://github.com/BioinfoMachineLearning/PoseBench.
@@ 1 Introduction
The field of drug discovery has long been challenged with a critical task: determining the structure of ligand molecules in complex with proteins and other key macromolecules (Warren et al., 2012). As accurately identifying such complex structures (in particular multi-ligand structures) can yield advanced insights into the binding dynamics and functional characteristics (and thereby, the medicinal potential) of numerous protein complexes in vivo, in recent years, significant resources have been spent developing new experimental and computational techniques for protein-ligand structure determination (Du et al., 2016). Over the last decade, machine learning (ML) methods for structure prediction have become indispensable components of modern structure determination at scale, with AlphaFold 2 for protein structure prediction being a recent hallmark example (Jumper et al., 2021).
@@
  • We introduce the first unified benchmark for protein-ligand structure generation that evaluates the performance of both recent DL-based methods as well as conventional methods for single and multi-ligand docking.
  • In contrast to several recent works on protein-ligand docking (Buttenschoen et al., 2024; Corso et al., 2024a), the benchmark results we present in this work are all within the context of apo (i.e., predicted) protein structures without known binding pockets, which notably enhances the practicality and real-world utility of this study’s findings.
  • Our newly proposed benchmark, PoseBench, enables specific insights into necessary areas of future work for accurate and generalizable protein-ligand structure generation, including that molecule pretraining seems to be key to generalizing to multi-ligand docking targets.
  • Our benchmark’s results also show that template-based algorithms for protein-ligand structure generation surpass the multi-ligand docking performance of several recent DL methods for protein-ligand docking, which suggests the importance of directly training and evaluating future DL methods on multi-ligand targets.
Figure 1: Overview of PoseBench, our comprehensive benchmark for practical ML modeling of single and multi-ligand protein complex structures in the context of apo (predicted) protein structures without known binding pockets (i.e., blind docking).
@@ 2 Related work
Benchmarking efforts for protein-ligand complexes. In response to the large number of new methods that have been developed for protein-ligand structure generation, recent works have introduced several new datasets and metrics with which to evaluate newly developed methods, with some of such benchmarking efforts focusing on modeling single-ligand protein interactions (Buttenschoen et al., 2024) and with others specializing in the assessment of multi-ligand protein interactions (Robin et al., 2023). One of the primary aims of this work is to bridge this gap by systematically assessing a selection of the latest (pocket-blind) structure generation methods within both interaction regimes in the context of unbound protein structures and ab initio complex structure prediction, efforts we describe in greater detail in the following section.
@@ 3.1 Preprocessed datasets
Astex Diverse dataset. The Astex Diverse dataset (Hartshorn et al., 2007) is a collection of 85 protein-ligand complexes composed of various drug-like molecules known to be of pharmaceutical or agrochemical interest, where a single representative ligand is present in each complex. This dataset can be considered an easy benchmarking set for many DL-based docking methods in that several of its proteins are known to overlap with the commonly used PDBBind (time-split) training dataset. Nonetheless, including this dataset for benchmarking allows one to determine the performance "upper bound" of each method’s docking capabilities for single-ligand protein complexes.
To perform apo docking with this dataset, we used ESMFold (Lin et al., 2023) to predict the complex structure of each of its proteins, where 5 of these 85 complexes were excluded from the effective benchmarking set due to being too large for structure prediction on an 80GB NVIDIA A100 GPU. For the remaining 80 complexes, we then optimally aligned their predicted protein structures to the corresponding ground-truth (holo) protein-ligand structures using the PLI-weighted root mean square deviation (RMSD) alignment algorithm originally proposed by Corso et al. (2022).
@@
Likewise to Astex Diverse, for the PoseBusters Benchmark set, we used ESMFold to predict the apo complex structures of each of its proteins. After filtering out 28 complexes for which the corresponding protein structure could not be predicted on an 80GB A100 GPU, we RMSD-aligned the remaining 280 predicted protein structures while optimally weighting each complex’s protein-ligand interface in the alignment. For the DockGen dataset, we refer readers to Appendix G.1.
CASP15 dataset. To assess the multi-ligand modeling capabilities of recent methods for protein-ligand structure generation, in this work, we introduce a curated version of the CASP15 PLI dataset introduced as a first-of-its-kind prediction category in the 15th Critical Assessment of Structure Prediction (CASP) competition (Robin et al., 2023) held in 2022. The CASP15 PLI set is originally comprised of 23 protein-ligand complexes, where we subsequently filter out 4 complexes based on (1) whether the CASP organizers ultimately assessed predictions for the complexes; (2) whether they are RNA-ligand complexes with no interacting protein chains; or (3) whether we could obtain a reasonably accurate prediction of the complex’s multimeric protein chains using either ESMFold or AlphaFold-based structure prediction on an 80GB A100 GPU (selecting for each complex the prediction which yielded the lowest-RMSD protein complex structure). Following this initial filtering step, we optimally align each remaining complex’s predicted protein structures to the corresponding ground-truth protein-(multi-)ligand structures, weighting each of the complex’s protein-ligand binding sites in the structural alignment.
@@
Sequence identity overlap. Note that for all four of the test datasets described above and listed in Table 1, we do not perform an analysis of the sequence identity overlap between these test datasets’ proteins and those of e.g., the PDBBind (time-split) training dataset, as (1) not all DL-based docking methods use PDBBind as their respective training datasets and (2) leaving the test complexes unfiltered according to sequence identity should, in principle, reflect many real-world use cases of these methods in which several (new) protein targets they are presented with may or may not be similar to what the methods have "seen" during training. Nevertheless, for an investigation of the sequence identity overlap between e.g., the PoseBusters Benchmark set and PDBBind, we refer interested readers to Buttenschoen et al. (2024). Furthermore, in Appendix F, we analyze the different types and frequencies of protein-ligand interactions natively found within the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively, to quantify the diversity of the (predicted) interactions each dataset can be used to evaluate.
@@ 4 Methods and experimental setup
As representative algorithms for conventional protein-ligand docking, we include AutoDock Vina (v1.2.5) (Trott and Olson, 2010) as well as a template-based modeling method for accurate ligand-protein complex structure prediction (TULIP) that we introduce in this work. To represent predictive ML docking algorithms, we include FABind (Pei et al., 2024) as well as the recently released version of RoseTTAFold 2 for all-atom structural modeling (i.e., RoseTTAFold-All-Atom) (Krishna et al., 2024). Lastly, for generative ML docking algorithms, we include DynamicBind (Lu et al., 2024), NeuralPLexer (Qiao et al., 2024), and the latest version of DiffDock referred to as DiffDock-L (Corso et al., 2024a) which is designed with binding site generalization as a key aim. Notably, AlphaFold 3 (Abramson et al., 2024) does not support generic SMILES string inputs, so we cannot benchmark it.
Additionally, we provide a method ensembling baseline (Ensemble) that uses (multi-)ligand structural consensus ranking (Con) (Roy et al., 2023) to rank its ligand structure predictions selected from the (intrinsically method-ranked) top-40 ligand conformations produced by each conventional, predictive, and generative ML method. This ensembling baseline is included to answer the question, "Which method produces the most consistent conformations in interaction with a protein complex?".
@@
[enumerate] \1 Formats for conventional methods are as follows: \2 Template-based methods such as TULIP are provided with an apo (predicted) protein structure and (fragment) ligand SMILES strings and are tasked with retrieving (PDB template (Bank, 1971)) ligand conformations residing in the same coordinate system as the given (predicted) protein structure following optimal molecular and structural alignment (Hu et al., 2018) with corresponding RDKit conformers of the input (query) ligand SMILES strings, where structural similarity with the query ligands is used to rank-order the selected (PDB template) conformations. \2 Molecular docking tools such as AutoDock Vina, which require specification of protein binding sites, are provided with not only a predicted protein structure but also the centroid coordinates of each (DiffDock-L-)predicted protein-ligand binding site residue. Such binding site residues are classified using a 4 Å protein-ligand heavy atom interaction threshold and using a 25 Å ligand-ligand heavy atom interaction threshold to define a "group" of ligands belonging to the same binding site and therefore residing in the same 25 Å3-sized binding site input voxel for AutoDock Vina. For interested readers, in Appendix G.1, we additionally report results using P2Rank (Krivák and Hoksza, 2018) to predict AutoDock Vina’s binding site centroid inputs.
\1
@@
Formats for predictive methods are as follows: \2 FABind is provided with a predicted protein structure as well as a ligand SMILES string, and it is then tasked with producing a (single) ligand conformation in complex with the given protein. \2 RoseTTAFold-All-Atom is provided with a (multi-chain) protein sequence as well as (fragment) ligand SMILES strings, and it is subsequently tasked with producing not only a (single) bound ligand conformation but also the bound protein conformation (as a representative ab initio structure generation method).
\1
@@
Formats for generative methods are as follows: \2 DiffDock-L is provided with a predicted protein structure and (fragment) ligand SMILES strings and is then tasked with producing (multiple rank-ordered) ligand conformations (for each fragment) for the given protein. Note that DiffDock-L does not natively support multi-ligand SMILES string inputs, so in this work, we propose a modified inference procedure for DiffDock-L which autoregressively presents each (fragment) ligand SMILES string to the model while providing the same predicted protein structure to the model in each inference iteration (reporting for each complex the average confidence score over all iterations). Notably, as an inference-time modification, this sampling formulation permits multi-ligand sampling yet cannot model multi-ligand interactions directly and therefore often produces ligand-ligand steric clashes. \2 As a single-ligand generative docking method, DynamicBind adopts the same input and output formats as DiffDock-L with the following exceptions: (1) the predicted input protein structure can be modified in response to (fragment) ligand docking; (2) the autoregressive inference procedure we adapted from that of DiffDock-L now provides DynamicBind with its own most recently generated protein structure in each (fragment) ligand inference iteration, thereby providing the model with partial multi-ligand interaction context; and (3) iteration-averaged confidence scores and predicted affinities are reported for each complex. Nonetheless, for both DiffDock-L and DynamicBind, such modified inference procedures highlight the importance in future work of retraining such generative methods directly on multi-ligand complexes to address such inference-time compromises. \2 Lastly, as a natively multi-ligand structure generation model pretrained using various 3D molecular and protein data sources, NeuralPLexer receives as its inputs a (multi-chain) protein sequence, a predicted protein (template) structure, as well as (fragment) ligand SMILES strings. The method is then tasked with producing (multiple rank-ordered) protein-ligand structure conformations for each input complex, using the method’s average predicted per-ligand heavy atom local Distance Difference Test (lDDT) score (Mariani et al., 2013) for rank-ordering.
@@
Prediction and evaluation procedures. Using the prediction formats above, the protein-ligand complex structures each method produces are subsequently evaluated using various structural accuracy and molecule validity metrics depending on whether the targets are single or multi-ligand complexes.
@@
Single-ligand evaluation. For single-ligand targets, we report each method’s percentage of (top-1) ligand conformations within 2 Å of the corresponding ground-truth ligand structure (RMSD $\leq$ 2 Å) as well as the percentage of such "correct" ligand conformations that are also considered to be chemically and structurally valid according to the PoseBusters software suite (Buttenschoen et al., 2024) (RMSD $\leq$ 2 Å & PB-Valid).
@@
Multi-ligand evaluation. Following CASP15’s official scoring procedure for protein-ligand complexes (Robin et al., 2023), for multi-ligand targets, we report each method’s percentage of "correct" (binding site-superimposed) ligand conformations (RMSD $\leq$ 2 Å) as well as violin plots of the RMSD and PLI-specific lDDT scores of its protein-ligand conformations across all (fragment) ligands within the benchmark’s multi-ligand complexes (see Appendix G for these plots). Notably, this final metric, referred to lDDT-PLI, allows one to evaluate specifically how well each method can model protein-ligand structural interfaces. We refer readings to Appendix D for formal definitions of these metrics. In the remainder of this work, we will discuss our benchmark’s results and their implications for the development of future complex structure generation methods.

### v4

@@ Abstract
The effects of ligand binding on protein structures and their in vivo functions carry numerous implications for modern biomedical research and biotechnology development efforts such as drug discovery. Although several deep learning (DL) methods and benchmarks designed for protein-ligand docking have recently been introduced, to date no prior works have systematically studied the behavior of docking methods within the broadly applicable context of (1) using predicted (apo) protein structures for docking (e.g., for applicability to unknown structures); (2) docking multiple ligands concurrently to a given target protein (e.g., for enzyme design); and (3) having no prior knowledge of binding pockets (e.g., for unknown pocket generalization). To enable a deeper understanding of docking methods’ real-world utility, we introduce PoseBench, the first comprehensive benchmark for broadly applicable protein-ligand docking. PoseBench enables researchers to rigorously and systematically evaluate DL docking methods for apo-to-holo protein-ligand docking and protein-ligand structure generation using both single and multi-ligand benchmark datasets, the latter of which we introduce for the first time to the DL community. Empirically, using PoseBench, we find that (1) DL methods consistently outperform conventional docking algorithms; (2) most recent DL docking methods fail to generalize to multi-ligand protein targets; and (3) training DL methods with physics-informed loss functions on diverse clusters of protein-ligand complexes is a promising direction for future work. Code, data, tutorials, and benchmark results are available at https://github.com/BioinfoMachineLearning/PoseBench.
@@ 1 Introduction
The field of drug discovery has long been challenged with a critical task: determining the structure of ligand molecules in complex with proteins and other key macromolecules (Warren et al., 2012). As accurately identifying such complex structures (in particular multi-ligand structures) can yield advanced insights into the binding dynamics and functional characteristics (and thereby, the medicinal potential) of numerous protein complexes in vivo, in recent years, significant resources have been spent developing new experimental and computational techniques for protein-ligand structure determination (Du et al., 2016). Over the last decade, machine learning (ML) methods for structure prediction have become indispensable components of modern structure determination at scale, with AlphaFold 2 for protein structure prediction being a hallmark example (Jumper et al., 2021).
@@
  • We introduce the first unified benchmark for protein-ligand structure generation that evaluates the performance of both recent DL-based methods as well as conventional methods for single and multi-ligand docking, which suggests that DL methods consistently outperform conventional docking techniques.
  • In contrast to several recent works on protein-ligand docking (Buttenschoen et al., 2024; Corso et al., 2024a), the benchmark results we present in this work are all within the context of high accuracy apo (i.e., AlphaFold 3-predicted) protein structures without known binding pockets, which notably enhances the broad applicability of this study’s findings.
  • Our newly proposed benchmark, PoseBench, enables specific insights into necessary areas of future work for accurate and generalizable protein-ligand structure generation, including that physics-informed inter-ligand clash losses seems to be key to generalizing to multi-ligand docking targets.
  • Our benchmark’s results also highlight the importance of considering rigorous (e.g., structure-based) dataset splits when training future DL docking methods and measuring their ability to recapitulate the ground-truth distributions of protein-ligand interactions within benchmark datasets.
Figure 1: Overview of PoseBench, our comprehensive benchmark for broadly applicable
ML modeling of single and multi-ligand protein complex structures in the context of apo (predicted) protein structures without known binding pockets (i.e., blind docking).
@@ 2 Related work
Benchmarking efforts for protein-ligand complexes. In response to the large number of new methods that have been developed for protein-ligand structure generation, recent works have introduced several new datasets and metrics with which to evaluate newly developed methods, with some of such benchmarking efforts focusing on modeling single-ligand protein interactions (Buttenschoen et al., 2024; Durairaj et al., 2024) and with others specializing in the assessment of multi-ligand protein interactions (Robin et al., 2023). One of the primary aims of this work is to bridge this gap by systematically assessing a selection of the latest (pocket-blind) structure generation methods within both interaction regimes in the context of unbound protein structures and ab initio complex structure prediction, efforts we describe in greater detail in the following section.
@@ 3.1 Preprocessed datasets
Astex Diverse dataset. The Astex Diverse dataset (Hartshorn et al., 2007) is a collection of 85 protein-ligand complexes composed of various drug-like molecules known to be of pharmaceutical or agrochemical interest, where a single representative ligand is present in each complex. This dataset can be considered an easy benchmarking set for many DL-based docking methods in that several of its proteins are known to overlap with the commonly used PDBBind (time-split) training dataset. Nonetheless, including this dataset for benchmarking allows one to determine the performance ”upper bound” of each method’s docking capabilities for single-ligand protein complexes.
To perform apo docking with this dataset, we used AlphaFold 3 (Abramson et al., 2024) to predict the complex structure of each of its proteins, where 5 of these 85 complexes were excluded from the effective benchmarking set due to being too large for docking with certain baseline methods on an 80GB NVIDIA A100 GPU. For the remaining 80 complexes, we then optimally aligned their predicted protein structures to the corresponding ground-truth (holo) protein-ligand structures using the PLI-weighted root mean square deviation (RMSD) alignment algorithm originally proposed by Corso et al. (2022).
@@
Likewise to Astex Diverse, for the PoseBusters Benchmark set, we used AlphaFold 3 to predict the apo complex structures of each of its proteins. After filtering out 28 complexes that certain baseline methods could not fit on an 80GB A100 GPU, we RMSD-aligned the remaining 280 predicted protein structures while optimally weighting each complex’s protein-ligand interface in the alignment. For the DockGen dataset, we refer readers to Appendix H.1.
CASP15 dataset. To assess the multi-ligand modeling capabilities of recent methods for protein-ligand structure generation, in this work, we introduce a curated version of the CASP15 PLI dataset introduced as a first-of-its-kind prediction category in the 15th Critical Assessment of Structure Prediction (CASP) competition (Robin et al., 2023) held in 2022. The CASP15 PLI set is originally comprised of 23 protein-ligand complexes, where we subsequently filter out 4 complexes based on (1) whether the CASP organizers ultimately assessed predictions for the complexes; (2) whether they are nucleic acid-ligand complexes with no interacting protein chains; or (3) whether we could obtain a reasonably accurate prediction of the complex’s multimeric protein chains using AlphaFold 3. Following this initial filtering step, we optimally align each remaining complex’s predicted protein structures to the corresponding ground-truth protein-(multi-)ligand structures, weighting each of the complex’s protein-ligand binding sites in the structural alignment.
@@
Dataset similarity analysis. For an investigation of the protein sequence similarity overlap between datasets such as the PoseBusters Benchmark set and the commonly-used PDBBind 2020 docking training dataset Liu et al. (2017), we refer interested readers to Buttenschoen et al. (2024). However, as a direct measure of the chemical and structural pocket similarity between PDBBind 2020 and the benchmark datasets employed in this work, in Appendix F.1, we analyze the different types and frequencies of protein-ligand pocket-level interactions natively found within the PDBBind 2020, Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively, to quantify the diversity of the (predicted) interactions each dataset can be used to evaluate and to obtain an estimate of the (pocket-based) generalization challenges posed by each dataset. In short, we find that the DockGen and CASP15 benchmark datasets are the most dissimilar compared to PDBBind 2020.
@@ 4 Methods and experimental setup
As representative algorithms for conventional protein-ligand docking, we include AutoDock Vina (v1.2.5) (Trott & Olson, 2010) as well as a template-based modeling method for ligand-protein complex structure prediction (TULIP) that we incorporate in this work to compare modern DL docking methods to the most common types of traditional docking algorithms (e.g., in the CASP15 competition (Xu et al., 2023)). For completeness, in Appendix G, we include a detailed description of the TULIP algorithm to provide interested readers with historical context regarding how such traditional docking techniques have typically been designed.
To represent predictive ML docking algorithms, we include FABind (Pei et al., 2024) as well as the recently released version of RoseTTAFold 2 for all-atom structural modeling (i.e., RoseTTAFold-All-Atom) (Krishna et al., 2024). Lastly, for generative ML docking algorithms, we include DynamicBind (Lu et al., 2024), NeuralPLexer (Qiao et al., 2024), Chai-1 (Chai-Discovery, 2024), and DiffDock-L (Corso et al., 2024a), the latest version of DiffDock, which is designed with pocket generalization as a key aim (n.b., through its use of ECOD (Cheng et al., 2014) structure-based cluster sampling). Notably, AlphaFold 3 (Abramson et al., 2024) does not currently support generic SMILES string inputs, so we cannot benchmark it.
Additionally, we provide a method ensembling baseline (Ensemble) that uses (multi-)ligand structural consensus ranking (Con) (Roy et al., 2023) to rank its ligand structure predictions selected from the (intrinsically method-ranked) top-3 ligand conformations produced by a subset of the DL baseline methods of this work (i.e., DiffDock-L, DynamicBind, NeuralPLexer, and RoseTTAFold-AA). This ensembling baseline is included to answer the question, ”Which of these DL methods produces the most consistent conformations in interaction with a protein complex?”.
@@
      Template-based (protein-fixed) methods such as TULIP are provided with an apo (predicted) protein structure and (fragment) ligand SMILES strings and are tasked with retrieving (PDB template (Bank, 1971)) ligand conformations residing in the same coordinate system as the given (predicted) protein structure following optimal molecular and structural alignment (Hu et al., 2018) with corresponding RDKit conformers of the input (query) ligand SMILES strings, where molecular similarity with the query ligands is used to rank-order the selected (PDB template) conformations.
    2. b)
      Molecular docking (protein-fixed) tools such as AutoDock Vina, which require specification of protein binding sites, are provided with not only a predicted protein structure but also the centroid coordinates of each (DiffDock-L-)predicted protein-ligand binding site residue. Such binding site residues are classified using a 4 Å protein-ligand heavy atom interaction threshold and using a 25 Å ligand-ligand heavy atom interaction threshold to define a ”group” of ligands belonging to the same binding site and therefore residing in the same 25 Å3-sized binding site input voxel for AutoDock Vina. For interested readers, for all four benchmark datasets, we also report results using P2Rank (Krivák & Hoksza, 2018) to predict AutoDock Vina’s binding site centroid inputs.
  2. 2.
    Formats for predictive methods are as follows:
    1. a)
@@
      FABind is provided with a predicted protein structure as well as a ligand SMILES string, and it is then tasked with producing a (single) ligand conformation in complex with the given (fixed-structure) protein.
    2. b)
      RoseTTAFold-All-Atom (AA) is provided with a (multi-chain) protein sequence as well as (fragment) ligand SMILES strings, and it is subsequently tasked with producing not only a (single) bound ligand conformation but also the bound (flexible) protein conformation (as a representative ab initio structure generation method).
  3. 3.
    Formats for generative methods are as follows:
    1. a)
@@
      DiffDock-L is provided with a predicted protein structure and (fragment) ligand SMILES strings and is then tasked with producing (multiple rank-ordered) ligand conformations (for each fragment) for the given (fixed-structure) protein. Note that DiffDock-L does not natively support multi-ligand SMILES string inputs, so in this work, we propose a modified inference procedure for DiffDock-L which autoregressively presents each (fragment) ligand SMILES string to the model while providing the same predicted protein structure to the model in each inference iteration (reporting for each complex the average confidence score over all iterations). Notably, as an inference-time modification, this sampling formulation permits multi-ligand sampling yet cannot model multi-ligand interactions directly and therefore often produces ligand-ligand steric clashes.
    2. b)
      As a single-ligand generative (flexible) docking method, DynamicBind adopts the same input and output formats as DiffDock-L with the following exceptions: (1) the predicted input protein structure is now flexible in response to (fragment) ligand docking; (2) the autoregressive inference procedure we adapted from that of DiffDock-L now provides DynamicBind with its own most recently generated protein structure in each (fragment) ligand inference iteration, thereby providing the model with partial multi-ligand interaction context; and (3) iteration-averaged confidence scores and predicted affinities are reported for each complex. Nonetheless, for both DiffDock-L and DynamicBind, such modified inference procedures highlight the importance in future work of retraining such generative methods directly on multi-ligand complexes to address such inference-time compromises.
    3. c)
      As a natively multi-ligand structure generation model trained using 3D molecular and protein data sources and a physics-informed (Van der Waals) clash loss, NeuralPLexer receives as its inputs a (multi-chain) protein sequence, a predicted protein (template) structure, as well as (fragment) ligand SMILES strings. The method is then tasked with producing multiple rank-ordered (flexible) protein-ligand structure conformations for each input complex, using the method’s average predicted per-ligand heavy atom local Distance Difference Test (lDDT) score (Mariani et al., 2013) for rank-ordering.
    4. d)
      Lastly, Chai-1 serves as a multi-ligand structure generation model (akin to AlphaFold 3) trained on diverse sequence-based PDB clusters and AlphaFold 2-predicted structures along with AlphaFold 3-based training losses. Following its default settings for inference, the model receives as its inputs a (multi-chain) protein sequence and (fragment) ligand SMILES strings, with no template structures or multiple sequence alignments provided. The method is then tasked with producing multiple rank-ordered (flexible) protein-ligand structure conformations for each input complex, using the method’s intrinsic ranking score (Abramson et al., 2024) for rank-ordering.
@@
Prediction and evaluation procedures. Using the prediction formats above, the protein-ligand complex structures each method produces are subsequently evaluated using various structural accuracy and molecule validity metrics depending on whether the targets are single or multi-ligand complexes. We refer readers to Appendix D for formal definitions of PoseBench’s structural metrics. Note that if a method’s prediction raises any errors in subsequent scoring stages (e.g., due to missing entities or formatting violations), the prediction is excluded from the evaluation.
@@
Single-ligand evaluation. For single-ligand targets, we report each method’s percentage of (top-1) ligand conformations within 2 Å of the corresponding ground-truth ligand structure (RMSD $\leq$ 2 Å) as well as the percentage of such ”correct” ligand conformations that are also considered to be chemically and structurally valid according to the PoseBusters software suite (Buttenschoen et al., 2024) (RMSD $\leq$ 2 Å & PB-Valid).
@@
Multi-ligand evaluation. Following CASP15’s official scoring procedure for protein-ligand complexes (Robin et al., 2023), for multi-ligand targets, we report each method’s percentage of ”correct” (binding site-superimposed) ligand conformations (RMSD $\leq$ 2 Å) as well as violin plots of the RMSD and PLI-specific lDDT scores of its protein-ligand conformations across all (fragment) ligands within the benchmark’s multi-ligand complexes (see Appendix H for these plots). Notably, this final metric, referred to lDDT-PLI, allows one to evaluate specifically how well each method can model protein-ligand structural interfaces. In the remainder of this work, we will discuss our benchmark’s results and their implications for the development of future structure generation methods.